# Fast Preemption: Forward-Backward Cascade Learning for Efficient and Transferable Proactive Adversarial Defense

### v6

@@ Abstract
Deep learning has advanced significantly but remains vulnerable to adversarial attacks, compromising its reliability. While conventional defenses typically mitigate perturbations post-attack, few studies explore proactive strategies for preemptive protection. This paper proposes Fast Preemption, a novel defense that neutralizes adversarial effects before third-party attacks. By employing distinct models for input labeling and feature extraction, Fast Preemption enables an efficient, transferable preemptive defense with state-of-the-art robustness across diverse systems. To further enhance efficiency, we introduce a forward-backward cascade learning algorithm that generates protective perturbations, leveraging forward propagation for rapid convergence and iterative backward propagation to mitigate overfitting. Executing in just three iterations, Fast Preemption surpasses existing training-time, test-time, and preemptive defenses. Additionally, we propose the first effective white-box adaptive reversion attack to evaluate the reversibility of preemptive defenses, demonstrating that our approach remains secure unless the backbone model, algorithm, and settings are entirely compromised. This work establishes a new paradigm for proactive adversarial defense.
@@ 1 Introduction
Deep learning models are highly susceptible to adversarial attacks [18], enabling adversaries to manipulate media such as images, audio, and video. For instance, an attacker can alter a facial image’s identity without perceptible changes, bypassing facial authentication systems [47], or introduce imperceptible noise to traffic signs, potentially compromising autonomous driving safety [15]. Defending against such attacks remains a critical yet unresolved challenge. Passive defenses [6, 29, 34, 22] enhance robustness at the cost of classification accuracy [54] and remain susceptible to adaptive attacks [7, 46]. Proactive defenses [32, 17], while mitigating threats preemptively, often lack real-time efficiency and require paired protection-inference modules, limiting transferability. Moreover, users relying on third-party models must contend with potential backdoor threats [39].
@@
Figure 1: A preemptive defense offers transferable protection against an adversarial attack by preemptively manipulating the image before it is attacked.
To address these challenges, we propose a proactive defense strategy, Fast Preemption, as illustrated in Fig. 1. Fast Preemption preemptively “attacks” media on the user side to enhance robustness before potential third-party attacks. Specifically, it learns robust features, predicts possible adversarial perturbations, and then amplifies these features while reversing the predicted perturbations for protection. As a result, any subsequent adversarial attack is inherently neutralized. This approach offers three key advantages. (i) Unlike passive defenses, Fast Preemption enhances threat resistance without compromising classification accuracy and requires no modifications to the classification system. (ii) It supports plug-and-play deployment without paired protection-inference modules, mitigating security and privacy concerns associated with external model providers [16, 39]. (iii) Traditional training- and test-time defenses, integrated within classification models, risk full-system compromise if breached. In contrast, Fast Preemption localizes vulnerabilities to individual users, as each employs distinct configurations. Thus, Fast Preemption satisfies four fundamental requirements:
  • Preserve clean accuracy and improve robust accuracy: Basic requirements for all defenses.
  • Not require ground truths: It is impractical to label each input manually.
  • Provide an efficient defense: Time cost per process must be satisfactory for user interaction.
  • Be transferable to other systems: Effect of protection must be transferable across various systems, so users do not have to use a paired module.
@@
We introduce Fast Preemption, a preemptive defense comprising three key components: a classifier model for automatic input labeling, a backbone model (distinct from the classifier) for feature learning, and the forward-backward (Fore-Back) cascade learning algorithm for efficiently computing transferable protective perturbations. We evaluate Fast Preemption on two datasets and thirteen models, defending against five representative adversarial attacks. Results demonstrate its superiority in clean and robust accuracy, transferability, and computational efficiency over benchmark training-time, test-time, and preemptive defenses [6, 29, 34, 32, 17]. For example, on CIFAR-10 [26], Fast Preemption outperformed the state-of-the-art (SOTA) preemptive defense [32], achieving 2.5% and 5.1% higher clean and robust accuracy, respectively, against AutoAttack (Standard) [11], while being 700 times faster.
Shannon famously stated, “One ought to design systems under the assumption that the enemy will immediately gain full familiarity with them” [43]. Following this principle, Bi-Level [32] first introduced a reversion attack but failed, asserting their defense was irreversible. In contrast, our reversion attack successfully circumvented their defense, making ours the first effective adaptive reversion attack in the white-box setting, termed Preemptive Reversion, to assess the reversibility of preemptive defenses. Preemptive Reversion applies a secondary preemptive defense to the already protected example and then reverses the newly introduced perturbations. Since the secondary defense generates perturbations similar to the first under identical conditions, this process effectively tests a defense’s resilience to reversion. Additionally, we propose a more rigorous evaluation protocol, which mandates a decrease in clean accuracy after preemptive defense (achieved by selecting a suboptimal classifier). If clean accuracy improves post-reversion, the attack is successful; if it further deteriorates due to distortion, reversion has failed.
Results of reversion attacks show that Preemptive Reversion effectively removes protective perturbations in a white-box setting, where all defense details are exposed. In contrast, blind JPEG distortion and adversarial purification [34] further degrade the image rather than neutralizing protections. However, this white-box assumption is unrealistic for real-world applications of Fast Preemption, as the framework allows users to customize their defense settings individually. Beyond the white-box scenario, existing reversion methods fail to circumvent Fast Preemption’s protection.
We summarize our contributions as follows: (i) We introduce Fast Preemption, a novel preemptive defense that enhances transferability without requiring ground truths or paired modules. It is the first approach to utilize distinct models for the classifier and backbone. (ii) We propose the forward-backward (Fore-Back) cascade learning algorithm, which accelerates preemptive defense, enabling efficient user interaction. (iii) We develop Preemptive Reversion, the first effective adaptive reversion attack for assessing the robustness of preemptive defenses in a white-box setting. Additionally, we establish a more comprehensive evaluation protocol to determine reversion success.
@@ 2 Related Works
Existing adversarial defenses can be classified into three main categories: training-time, test-time, and preemptive defenses. A detailed review is provided in Appendix F.
@@
Test-time defenses are typically divided into detection and purification. Detection models identify AEs by distinguishing them from benign inputs [9, 56, 27, 22], but detection is challenging for unseen attacks, as detectors are often trained on known AEs. Furthermore, detection may misclassify benign images as AEs, reducing clean accuracy. In contrast, adversarial purification aims to remove perturbations in AEs. Early approaches used image distortion for purification [48, 20, 37], while recent methods generate benign replacements regardless of input type [38, 34]. A key limitation of purification is that the difference between original and generated examples can degrade clean accuracy, and counterattack potential remains [7, 46].
@@
Alfarra et al. [3] proposed a test-time defense incorporating preemptive protection as an anti-adversary layer guided by a standard classifier. However, their method is inherently vulnerable if the classifier is compromised, a limitation they explicitly acknowledge. This weakness undermines its practical viability in test-time scenarios. While they describe this setting as the “least realistic,” we argue that it aligns with the standard white-box attack model, where adversaries have full access to the classifier. Under even stricter assumptions, adversaries may also be aware of the defense mechanism itself (white-box reversion). Moreover, their approach relies solely on forward propagation learning, increasing susceptibility to overfitting.
@@ 3.2 Framework
Classifier: Correct input labels determines the correct optimization direction for defense. However, for efficient user interaction, it is impractical to label each input manually. Classifier $f_{c}$ labels these inputs automatically. Therefore, $y$ in Eq. 3 and Eq. 4 can be replaced by $f_{c}(x)$. The framework is integrated with user application before being attacked, so $f_{c}$ does not have to show any transferability or resistance against threats, but it must offer the best classification accuracy. Processing images with ground truths can be regarded as using a perfect classifier.
@@ 3.3 Forward-Backward Cascade Learning
The Fore-Back cascade learning algorithm is used to compute protective perturbations. The input is first moved away from the classification boundaries of all incorrect classes so that the manipulated input will be easier to be classified as the ground truth. We define this as forward propagation ($\rightarrow$). Forward propagation can quickly attain convergence (less time cost) but causes overfitting on the backbone model so that the defense is less transferable to other victim models. Backward propagation optimization ($\leftarrow$) is thus conducted following forward propagation. Backward propagation reverses the adversarial perturbations to make them protective perturbations, which effectively alleviates overfitting and thereby improves transferability. In summary, the forward and backward propagation both enhance robustness, but in different optimization directions. The complete algorithm is presented in Algorithm 1 of Appendix D.
@@ 3.3.3 Combining Forward and Backward Propagation
where $x_{*}$ denotes $x_{t,i+1}$ or $x_{t+1}$. Equation 12 indicates that both forward and backward propagation make inputs more robust but with different learning directions for loss $\ell$.
@@ 4 Adaptive Attack: Preemptive Reversion
A reversion attack is used to assess the reversibility of a preemptive defense. While Bi-Level [32] first proposed a reversion attack but failed to reverse their defense, claiming it was irreversible, our reversion attack successfully overcame their defense, making ours the first effective reversion attack. In this section, we present the algorithm for our proposed Preemptive Reversion and a more comprehensive protocol for evaluating the effectiveness of reversion attacks. Preemptive Reversion is the only viable approach to reverse preemptive defenses in the white-box scenario. However, without full knowledge of the backbone model and settings, our Fast Preemption defense remains difficult to reverse, ensuring its reliability.
@@ 4.1 Algorithm
When running preemptive defenses sequentially for the same input and settings, the perturbations are mostly the same for two rounds (see Fig. 5). On the basis of this finding, we devised the first effective white-box adaptive reversion attack, i.e., Preemptive Reversion, as illustrated in Fig. 3. The complete algorithm is presented in Algorithm 2 of Appendix D. Using RE $x^{r}$ generated by a preemptive defense method, the Preemptive Reversion algorithm first runs a secondary preemptive defense upon $x^{r}$ to obtain $(x^{r})^{r}$. Then, the Preemptive Reversion algorithm reverses the secondary perturbations from $x^{r}$ to obtain $\tilde{x}$, which is the prediction of the original input $x$:
@@ 4.2 Attack Scenarios
  • White-Box: The attacker has complete knowledge of both the backbone model and defense algorithm (including settings), enabling them to replicate the secondary defense. This scenario, while the strongest, is impractical due to varying user settings.
@@
Our proposed Preemptive Reversion attack is the first effective method for evaluating the reversibility of preemptive defenses, including Fast Preemption, but it is only feasible in the white-box scenario. Without full knowledge of the defense, Fast Preemption remains reliable.
@@ 5 Experiment Settings
Datasets, Classifier, and Backbone Models. We conducted experiments on two datasets: CIFAR-10 [26] and ImageNet [2]. For CIFAR-10, ResNeXt29-32x4d [21] and PreActResNet-18 [19] were selected as the classifier and backbone model, respectively. For ImageNet, DeiT Base [45] and ConvNeXt-L+ConvStem [44] were selected as the classifier and backbone model, respectively. The model architectures are listed in Tab. 9 of Appendix E.
Adversarial attacks. We evaluated defenses against five adversarial attacks: white-box PGD [31], AutoAttack [11], and Carlini-Wagner (CW) attack [7]; black-box Deformation-Constrained Warping (DeCoW) [28] and Meta Conditional Generator (MCG) [53]. For AutoAttack, we used both its “Standard” version (combining APGD-CE, APGD-DLR, FAB [10], and Square Attack [4]) and its “Random” version (combining APGD-CE and APGD-DLR, with expectation over time (EOT) equal to 20 [5]).
Reversion attacks. We examined: (i) whether protected samples induced by preemptive defenses cannot be easily reversed to their original unprotected state, which we define as “irreversible”; and (ii) the effectiveness of the proposed adaptive reversion attack, Preemptive Reversion. Along with Preemptive Reversion, we also ran reversion attacks using blind distortion (JPEG compression) and adversarial purification (DiffPure [34]).
Training-time defenses. We compared Fast Preemption with WideResNet-94-16 [6] and Swin-L [29] on the CIFAR-10 and ImageNet datasets, respectively. The architectures of these models differ from the backbone models. To assess the transferability of preemptive defenses, we conducted experiments across additional nine victim models, with results presented in Tab. 7 of Appendix C.
Test-time defenses. We used DiffPure [34] as the benchmark test-time defense. The inputs were purified by DiffPure and then fed into WideResNet-94-16 [6] or Swin-L [29] depending on the dataset.
Preemptive defenses. We compared Fast Preemption with the SOTA preemptive defense, Bi-Level [32], and the latest preemptive defense, $A^{5}$ [17].
5.1 Evaluation metrics
We evaluated performance using clean accuracy and robust accuracy, where higher values indicate better defense effectiveness. Clean accuracy was computed for benign images (original and REs), while robust accuracy was computed for AEs. Additionally, to assess image quality, we calculated the structural similarity index measure (SSIM) [49] and Learned Perceptual Image Patch Similarity (LPIPS) [55] relative to the original images. Higher SSIM and lower LPIPS values indicate subtler changes in the images.
@@ 5.1 Evaluation metrics
Fast Preemption (ours) + DiffPure [34] 0.34 94.2 89.1 92.8 89.7
@@
Fast Preemption (ours) + DiffPure [34] 2.5 93.3 82.9 87.6 83.4

### v7

# Fast Preemption: Forward-Backward Cascade Learning for Efficient and Transferable Preemptive Adversarial Defense

@@ Abstract
Deep learning has made significant strides but remains susceptible to adversarial attacks, undermining its reliability. Most existing research addresses these threats after attacks happen. A growing direction explores preemptive defenses like mitigating adversarial threats proactively, offering improved robustness but at cost of efficiency and transferability. This paper introduces Fast Preemption, a novel preemptive adversarial defense that overcomes efficiency challenges while achieving state-of-the-art robustness and transferability, requiring no prior knowledge of attacks and target models. We propose a forward-backward cascade learning algorithm, which generates protective perturbations by combining forward propagation for rapid convergence with iterative backward propagation to prevent overfitting. Executing in just three iterations, Fast Preemption outperforms existing training-time, test-time, and preemptive defenses. Additionally, we introduce an adaptive reversion attack to assess the reliability of preemptive defenses, demonstrating that our approach remains secure in realistic attack scenarios.
@@ 1 Introduction
Deep learning models are highly susceptible to adversarial attacks [18], enabling adversaries to manipulate media such as images, audio, and video. For instance, an attacker can alter a facial image’s identity without perceptible changes, bypassing facial authentication systems [44], or introduce imperceptible noise to traffic signs, potentially compromising autonomous driving safety [15]. Defending against such attacks remains a critical yet unresolved challenge.
Training-time and test-time defenses [6, 26, 31, 22] mitigate adversarial threats after attacks happen, improving robustness but reducing classification accuracy [51] and remaining vulnerable to adaptive attacks [7, 43]. In contrast, preemptive defenses [29, 17] preemptively modify media to neutralize future adversarial manipulations. However, preemptive defenses suffer from inefficiency and rely on paired protection-inference modules, limiting transferability. Additionally, the protection modules in preemptive defenses are dependent on the paired, provided models, rendering them susceptible to potential backdoor threats [36].
@@
Figure 1: The proposed Fast Preemption offers transferable protection against an adversarial attack by preemptively manipulating the image before it is attacked.
To address these challenges, we propose Fast Preemption, a novel preemptive defense illustrated in Fig. 1. Fast Preemption learns robust features, predicts potential adversarial perturbations, and amplifies these features while reversing the predicted perturbations, effectively neutralizing subsequent attacks. It offers six key advantages over other defenses:
  • No reliance on ground truths: While ground-truth labels offer the highest protection, manual labeling is less practical. Fast Preemption operates without this requirement.
  • Plug-and-play: Unlike training-time defenses, Fast Preemption requires no modifications to the classification system. Unlike existing preemptive defenses, it eliminates reliance on paired protection-inference modules, reducing security and privacy risks associated with external model providers [16, 36].
  • Transferable: Fast Preemption provides generalized protection without requiring prior knowledge of target models or attack algorithms.
  • Efficient: A novel learning algorithm greatly improves efficiency compared to existing preemptive defenses.
  • State-of-the-art (SOTA) clean and robust accuracy: Unlike training-time and test-time defenses, Fast Preemption enhances threat resistance without compromising classification accuracy. It achieves the best clean and robust accuracy among existing adversarial defenses.
  • Resilient to adaptive attacks: Fast Preemption allows users to set protection strength and select a backbone model independently of the target model, enhancing resilience to adaptive attacks beyond existing test-time and preemptive defenses.
To achieve these objectives, we design Fast Preemption with three key components: a user-configurable classifier for automatic input labeling, a user-configurable backbone model (which may differ from the classifier) for feature learning, and a forward-backward (Fore-Back) cascade learning algorithm for efficiently computing transferable protective perturbations. The Fore-Back cascade learning algorithm integrates forward and backward propagation to overcome their individual limitations. Forward propagation ensures rapid convergence but overfits to the backbone model, reducing transferability. Conversely, backward propagation mitigates overfitting but requires many iterations, making it computationally expensive. By combining both, Fore-Back achieves efficient and robust protection.
@@
We evaluate Fast Preemption on two datasets and thirteen models, defending against four representative adversarial attacks. Results demonstrate its superiority in clean and robust accuracy, transferability, and computational efficiency over benchmark training-time, test-time, and preemptive defenses. For example, on CIFAR-10 [23], Fast Preemption outperforms the SOTA preemptive defense [29], achieving 2.5% and 5.1% higher clean and robust accuracy, respectively, against AutoAttack (Standard) [11], while being 700 times faster.
We further explore stronger adaptive attack scenarios where the attacker has full knowledge of our defense and attempts to reverse its protective perturbations. Moon et al. [29] introduced the first reversion attack but failed, claiming their preemptive defense was irreversible. We propose Preemptive Reversion, a more powerful reversion attack that successfully bypasses their defense in a white-box setting (i.e., both the defense algorithms and settings have been compromised). Preemptive Reversion applies a secondary preemptive defense to an already protected example and then reverses the newly introduced perturbations. Since the secondary defense generates perturbations similar to the first under identical conditions, this process effectively evaluates a defense’s susceptibility to reversion.
We employ Preemptive Reversion alongside methods such as blind JPEG distortion and adversarial purification [31] to assess the reliability of Fast Preemption against adaptive attacks. Since Fast Preemption allows for user-configurable parameterization (e.g., selection of the backbone model), and these configurations remain confidential, all adaptive attacks fail to bypass its protection, highlighting its robustness. However, in the less-realistic scenario where the user configurations are compromised, only Preemptive Reversion successfully removes the protective perturbations without introducing significant distortion, demonstrating the aggressiveness of the proposed attack. To our knowledge, Preemptive Reversion remains the only viable approach for evaluating the reliability of preemptive defenses against adaptive attacks.
We summarize our contributions as follows: (i) We introduce Fast Preemption, a more effective plug-and-play preemptive defense that requires neither ground truths nor prior knowledge of attacks or target models, while also providing resistance to adaptive attacks. (ii) We propose the Fore-Back cascade learning algorithm, which significantly enhances the transferability and learning speed of preemptive defenses. (iii) We develop Preemptive Reversion, a more robust tool for evaluating the reliability of preemptive defenses against adaptive attacks.
@@ 2 Related Works
Existing adversarial defenses can be classified into three main categories: training-time, test-time, and preemptive defenses.
@@
Test-time defenses are divided into confusion, detection, and purification. Confusion methods integrate multiple layers or models, so that if an attacker bypasses one, other layers can resist the threat [42, 3]. However, Alfarra et al. [42] and Wang et al. [44] acknowledge that this strategy is vulnerable when the attacker is aware of the defense and targets all layers in a single attack. Detection models identify AEs by distinguishing them from benign inputs [9, 53, 24, 22], but detection is challenging for unseen attacks, as detectors are often trained on known AEs. Furthermore, detection may misclassify benign images as AEs, reducing clean accuracy. In contrast, adversarial purification aims to remove perturbations in AEs. Early approaches used image distortion for purification [45, 20, 34], while recent methods generate benign replacements regardless of input type [35, 31]. A key limitation of purification is that the difference between original and generated examples can degrade clean accuracy, and counterattack potential remains [7, 43].
@@ 3.2 Framework
Classifier: Correct input labels determines the correct optimization direction for defense. However, for efficient user interaction, it is less practical to label each input manually. Classifier $f_{c}$ labels these inputs automatically. Therefore, $y$ in Eq. 3 and Eq. 4 can be replaced by $f_{c}(x)$. The framework is integrated with user application before being attacked, so $f_{c}$ does not have to show any transferability or resistance against threats, but it must offer the best classification accuracy. Processing images with ground truths can be regarded as using a perfect classifier.
@@ 3.3 Forward-Backward Cascade Learning
The Fore-Back cascade learning algorithm is used to compute protective perturbations. The input is first moved away from the classification boundaries of all incorrect classes so that the manipulated input will be easier to be classified as the ground truth. We define this as forward propagation ($\rightarrow$). Forward propagation can quickly attain convergence (less time cost) but causes overfitting on the backbone model so that the defense is less transferable to other victim models. Backward propagation optimization ($\leftarrow$) is thus conducted following forward propagation. Backward propagation reverses the adversarial perturbations to make them protective perturbations, which effectively alleviates overfitting and thereby improves transferability. In summary, the forward and backward propagation both enhance robustness, but in different optimization directions. The complete algorithm is presented in Algorithm 1 of Appendix A.
@@ 3.3.3 Combining Forward and Backward Propagation
where $x_{*}$ denotes $x_{t,i+1}$ or $x_{t+1}$. Eq. 12 indicates that both forward and backward propagation make inputs more robust but with different learning directions for loss $\ell$.
@@ 4 Adaptive Attack: Preemptive Reversion
A reversion attack is used to assess the reliability of a preemptive defense. While Moon et al. [29] first proposed a reversion attack against their proposed defense, Bi-Level, but failed to reverse it, claiming it was irreversible, we propose a stronger reversion attack successfully overcame Bi-Level. In this section, we present the algorithm for our proposed Preemptive Reversion and a more comprehensive protocol for evaluating the effectiveness of reversion attacks.
@@ 4.1 Algorithm
When running preemptive defenses sequentially for the same input and settings, the perturbations are mostly the same for two rounds (see Fig. 5). On the basis of this finding, we devised the adaptive reversion attack, i.e., Preemptive Reversion, as illustrated in Fig. 3. The complete algorithm is presented in Algorithm 2 of Appendix A. Using RE $x^{r}$ generated by a preemptive defense method, the Preemptive Reversion algorithm first runs a secondary preemptive defense upon $x^{r}$ to obtain $(x^{r})^{r}$. Then, the Preemptive Reversion algorithm reverses the secondary perturbations from $x^{r}$ to obtain $\tilde{x}$, which is the prediction of the original input $x$:
@@ 4.2 Attack Scenarios
  • White-Box: The attacker has complete knowledge of both the backbone model and defense algorithm (including settings), enabling them to replicate the secondary defense. This scenario, while the strongest, is less realistic due to varying user settings.
@@
Our proposed Preemptive Reversion attack is effective for evaluating the reversibility of preemptive defenses, including Fast Preemption, but it is only feasible in the white-box scenario. Without full knowledge of the defense, Fast Preemption remains reliable.
@@ 5 Experiment Settings
Datasets, Classifier, and Backbone Models. We conducted experiments on two datasets: CIFAR-10 [23] and ImageNet [2]. For CIFAR-10, ResNeXt29-32x4d [21] and PreActResNet-18 [19] were selected as the classifier and backbone model, respectively. For ImageNet, DeiT Base [41] and ConvNeXt-L+ConvStem [40] were selected as the classifier and backbone model, respectively. The model architectures are listed in Tab. 5 of Appendix B.
Adversarial attacks. We evaluated defenses against four adversarial attacks: white-box attacks include AutoAttack [11] and Carlini-Wagner (CW) [7]; black-box attacks include Deformation-Constrained Warping (DeCoW) [25] and Meta Conditional Generator (MCG) [50]. For AutoAttack, we used both its “Standard” version (combining APGD-CE, APGD-DLR, FAB [10], and Square Attack [4]) and its “Random” version (combining APGD-CE and APGD-DLR, with expectation over time (EOT) equal to 20 [5]). Reversion attacks. We examined: (i) whether protected samples induced by preemptive defenses cannot be easily reversed to their original unprotected state, which we define as “irreversible”; and (ii) the effectiveness of the proposed adaptive reversion attack, Preemptive Reversion. Along with Preemptive Reversion, we also ran reversion attacks using blind distortion (JPEG compression) and adversarial purification (DiffPure [31]).
Training-time defenses. We compared Fast Preemption with WideResNet-94-16 [6] and Swin-L [26] on the CIFAR-10 and ImageNet datasets, respectively. The architectures of these models differ from the backbone models. To assess the transferability of preemptive defenses, we conducted experiments across additional nine victim models, with results presented in Tab. 6 of Appendix C. Test-time defenses. We used DiffPure [31] as the benchmark test-time defense. The inputs were purified by DiffPure and then fed into WideResNet-94-16 [6] or Swin-L [26] depending on the dataset. Preemptive defenses. We compared Fast Preemption with the SOTA preemptive defense, Bi-Level [29], and the latest preemptive defense, $A^{5}$ [17].
Evaluation metrics. We evaluated performance using clean accuracy and robust accuracy, where higher values indicate better defense effectiveness. Clean accuracy was computed for benign images (original and REs), while robust accuracy was computed for AEs. Additionally, to assess image quality, we calculated the structural similarity index measure (SSIM) [46] and Learned Perceptual Image Patch Similarity (LPIPS) [52] relative to the original images. Higher SSIM and lower LPIPS values indicate subtler changes in the images.
@@
Combination Ours+[31] 0.34 94.2 89.1 92.8 89.7
@@
Combination Ours+[31] 2.5 93.3 82.9 87.6 83.4